OpenStudy (anonymous):

Simplify (3x^-3/4x^3)^-2

OpenStudy (anonymous):

\[\frac{3x^{-3}}{4x^{3}}^{-2}\] \[\frac{3}{4x^{3}x^{3}}^{-2}\] \[\frac{3}{4x^{6}}^{-2}\] \[\sqrt{3x^6/4}\] \[\frac{\sqrt{3}x^3}{2}\]

OpenStudy (anonymous):

\[\text{Nice}\]